# CKnowEdit: A New Chinese Knowledge Editing Dataset for Linguistics, Facts, and Logic Error Correction in LLMs

### v3

@@ Abstract
Jizhan Fang1,thanks: Equal contribution and shared co-first authorship.     Tianhe Lu1,11footnotemark: 1     Yunzhi Yao1,11footnotemark: 1 Ziyan Jiang1     Xin Xu2     Ningyu Zhang1,thanks: Corresponding Author.     Huajun Chen1,22footnotemark: 2 1Zhejiang University        2University of California, San Diego {fangjizhan, yyztodd, zhangningyu}@zju.edu.cnxinxucs@ucsd.edu
@@ 1 Introduction
The reliance on static training data and the lack of explicit knowledge representation in Large Language Models often lead to issues such as hallucinations, bias, and offensive outputs Zhao et al. (2023); Huang et al. (2023a); Liu et al. (2023); Sun et al. (2024b). These limitations become particularly pronounced when LLMs operate in complex domains or languages, such as Chinese. As shown in Figure 1, Chinese is a highly complex and linguistically unique system and presents three distinct challenges for LLMs versus Indo-European languages Luelsdorff (1994); Matthiessen (2023); Xu et al. (2023): $(i)$ Linguistic Complexity: Characters intricately blend shape, sound, and meaning through composition and contextual pronunciation shifts, while flexible grammar and cultural elements (poetry, idioms, etc.) evolved over millennia. $(ii)$ Culture-Laden Facts: Untranslatable contexts in specific facts like geographical/historical terms. $(iii)$ Language-Specific Logic: Context-dependent reasoning patterns that rely on implicit connectors and topic prominence over subject-predicate structures, often leading to misalignment in logical chain extraction.
@@ 4.1 Experiment Settings
Figure 4: Main results. We do not report the locality of Ancient Poetry, Proverbs, Idioms and Facts Knowledge because it is challenging to find out-scope knowledge that is both relevant to and distinct from the target knowledge when we construct the locality field.
@@
Unlike conventional knowledge editing evaluations that use token/logit-level metrics with teacher-forcing automation Yao et al. (2023), we adopt open-ended text generation to assess edited models in more practical and challenging scenarios. While some studies Deng et al. (2024) under similar setups use metrics like ROUGE-L or semantic similarity, we find these metrics often fail to reflect true text quality. For instance, ROUGE-L is heavily skewed by text length—shorter reference texts paired with longer model outputs lead to unreliable scores.
@@ 4.4 Human Evaluation
To verify the effectiveness of our designed automatic GPT-4 score for CKnowEdit evaluation, we randomly select 70 data samples from all knowledge types, along with outputs from 4 baseline models for human evaluation by our contracted annotators. From the human evaluation results, the overall correlation coefficient across all 4 metrics between the automatic and human evaluation is 0.70, indicating a high consistency between GPT-4 scores and human preferences.
@@ 6 Conclusion
Since the original intention of this work is to study knowledge with distinctive Chinese linguistic characteristics, and Chinese linguistic knowledge and Chinese logical knowledge better reflect these features, the quantity of these two types of knowledge in CKnowEdit is significantly greater than that of factual knowledge. This is a limitation of CKnowEdit. In addition, using GPT-4 to evaluate the output of other LLMs is already a widely used method. Although using GPT-4 to evaluate GPT-4 may be biased, using GPT-4 to evaluate other models still has reference value. In addition, not only for the tasks we proposed, but also for other tasks, the community is still actively exploring how to effectively evaluate LLMs. As a temporary compromise, we remind readers that they should interpret the GPT-4 scores carefully.
@@ Appendix A Data Source Website
Table 1: Results (Edit Success / Generalization / Portability / Locality) of pre-edit and post-edit with 4 knowledge editing methods for LLMs. We color the metrics where the editing results in negative gains red. The underlined numbers correspond to the best method for every metric.

### v4

@@ Abstract
Jizhan Fang1thanks: Equal contribution and shared co-first authorship.     Tianhe Lu1,11footnotemark: 1     Yunzhi Yao1 Ziyan Jiang1     Xin Xu2     Huajun Chen1    Ningyu Zhang1thanks: Corresponding Author. 1Zhejiang University        2University of California, San Diego {fangjizhan, yyztodd, zhangningyu}@zju.edu.cnxinxucs@ucsd.edu
@@ 1 Introduction
The reliance on static training data and the lack of explicit knowledge representation in Large Language Models often lead to issues such as hallucinations, bias, and offensive outputs Zhao et al. (2023); Huang et al. (2023a); Liu et al. (2023); Sun et al. (2024b); Chen (2024). These limitations become particularly pronounced when LLMs operate in complex domains or languages, such as Chinese. As shown in Figure 1, Chinese is a highly complex and linguistically unique system and presents three distinct challenges for LLMs versus Indo-European languages Luelsdorff (1994); Matthiessen (2023); Xu et al. (2023); Wen et al. (2023): $(i)$ Linguistic Complexity: Characters intricately blend shape, sound, and meaning through composition and contextual pronunciation shifts, while flexible grammar and cultural elements (poetry, idioms, etc.) evolved over millennia. $(ii)$ Culture-Laden Facts: Untranslatable contexts in specific facts like geographical/historical terms. $(iii)$ Language-Specific Logic: Context-dependent reasoning patterns that rely on implicit connectors and topic prominence over subject-predicate structures, often leading to misalignment in logical chain extraction.
@@ 3.4 Quality Assurance
3.5 Field and Usage Specifications
As shown in Figure 7, ’prompt’ and ’target new’ fields constitute the target knowledge we aim to edit. ’target old’ represents the original incorrect response generated by Qwen-7B-Chat during our data filtering phase.
With the continuous advancement of LLMs, their capabilities are growing increasingly powerful. As a result, some data in CKnowEdit may already be correctly answered by more capable models. However, this does not diminish the significance of CKnowEdit for editing purposes. During usage, ’target new’ and ’target old’ fields can be swapped to construct counterfactual data (as is commonly done in many editing datasets). This approach not only preserves the essence of the editing operation but also maintains the unique Chinese characteristics of this dataset.
@@ 4.1 Experiment Settings
Figure 4: Main results. We do not report the locality of Ancient Poetry, Proverbs, Idioms and Facts Knowledge because it is challenging to find out-scope knowledge that is both relevant to and distinct from the target knowledge when we construct the locality field. The corresponding numerical results are presented in the Table1.
@@
Unlike conventional knowledge editing evaluations that use token/logit-level metrics with teacher-forcing automation, we adopt open-ended text generation to assess edited models in more practical and challenging scenarios. While some studies Deng et al. (2024) under similar setups use metrics like ROUGE-L or semantic similarity, we find these metrics often fail to reflect true text quality. For instance, ROUGE-L is heavily skewed by text length—shorter reference texts paired with longer model outputs lead to unreliable scores. A recent study  Yang et al. (2025) has also revealed specific inadequacies in traditional evaluation methods.
@@ 4.4 Human Evaluation
To verify the effectiveness of our designed automatic GPT-4 score for CKnowEdit evaluation, from the outputs of 4 baseline models edited by 5 different methods (totaling 20 output categories), we select 70 samples per category for human evaluation conducted by our 5 contracted annotators under rigorous assessment standards. From the human evaluation results, the overall correlation coefficient across all 4 metrics between the automatic and human evaluation is 0.70, indicating a high consistency between GPT-4 scores and human preferences.
@@ 6 Conclusion
Imbalanced Distribution of Data Types
Since the original intention of CKnowEdit is to study knowledge with distinctive Chinese linguistic characteristics, and Chinese linguistic knowledge or logical knowledge better reflect these features, the quantity of these two types of knowledge in CKnowEdit is significantly greater than that of factual knowledge.
Experimental Setup
Due to computational resource constraints, all experiments in this study were conducted solely under the single-edit setting, without investigating batch edit or sequential edit scenarios.
LLM-as-a-judge
The use of GPT-4 as an evaluator for other LLMs has become a widely adopted practice in the field. While employing GPT-4 to evaluate itself may introduce inherent biases, its assessments of other models still provide valuable reference points. It’s important to note that the challenge of effectively evaluating LLMs extends beyond our specific tasks - the research community continues to actively investigate robust evaluation methodologies across various applications. As an interim solution, we advise readers to interpret GPT-4 evaluation scores with appropriate caution, recognizing both their utility and limitations.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Yongjiang Talent Introduction Programme (2021A-156-G), Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003), Ningbo Natural Science Foundation (2024J020), Information Technology Center and State Key Lab of CAD&CG, Zhejiang University. We gratefully acknowledge the support of Zhejiang University Education Foundation Qizhen Scholar Foundation.
@@ Appendix A Data Source Website
Table 1: Numerical results (Edit Success / Generalization / Portability / Locality) of pre-edit and post-edit with 5 knowledge editing methods for 4 baseline LLMs.